How much water must you add to 250.0mL of orange juice concentrate if you want the final juice to be 15.0% of the strength of the original?

Found 2 solutions by stanbon, rwm:
Answer by stanbon(75887)

How much water must you add to 250.0mL of orange juice concentrate if you want the final juice to be 15.0% of the strength of the original?
-----------------
Equation:
OJ + OJ = OJ
1.00*250 + 0*x = 0.15(250+x)
---
Multiply thru by 100 to get:
100*250 = 15*250 + 15x
---
15x = 85*250
---
x = 1416.67 mL (amt of water needed)
--------------------------------------
Cheers,
Stan H.

Answer by rwm(914)

so 250 ml will be 15% of the total juice mix
.15x=250
x will be the total mixture.
1666.67
so subtract 250 from the total to get the water added
1666.67-250=1416.67 ml of water